A boat is sailing 6.17 m/s directly cast across the harbor when a gust of wind comes up, the boat changes its direction to 11.0°, north of cast and its speed increases to 7.2 m/s. What is the change in velocity of the boat?
1
Expert's answer
2020-02-03T14:00:30-0500
"\Delta \vec{v}=(7.2\cos{11},7.2\sin{11})-(6.17,0)"

"\Delta \vec{v}=(0.898,1.374)\frac{m}{s}"

"\Delta v=\sqrt{0.898^2+1.374^2}=1.64\frac{m}{s}"

Direction:


"\theta=\arctan{\frac{1.374}{0.898}}=56.8\degree N\ of\ E"